# PRISMM-Bench: A Benchmark of Peer-Review Grounded Multimodal Inconsistencies

### v2

@@ Abstract
Large Multimodal Models (LMMs) are increasingly applied to scientific research, yet it remains unclear whether they can reliably understand and reason over the multimodal complexity of papers. A central challenge lies in detecting and resolving inconsistencies across text, figures, tables, and equations, issues that are often subtle, domain-specific, and ultimately undermine clarity, reproducibility, and trust. Existing benchmarks overlook this issue, either isolating single modalities or relying on synthetic errors that fail to capture real-world complexity. We introduce PRISMM-Bench (Peer-Review-sourced Inconsistency Set for Multimodal Models), the first benchmark grounded in real reviewer-flagged inconsistencies in scientific papers. Through a multi-stage pipeline of review mining, LLM-assisted filtering and human verification, we curate 262 inconsistencies from 242 papers. Based on this set, we design three tasks, namely inconsistency identification, remedy and pair matching, which assess a model’s capacity to detect, correct, and reason over inconsistencies across different modalities. Furthermore, to address the notorious problem of choice-only shortcuts in multiple-choice evaluation, where models exploit answer patterns without truly understanding the question, we further introduce structured JSON-based answer representations that minimize linguistic biases by reducing reliance on superficial stylistic cues. We benchmark 21 leading LMMs, including large open-weight models (GLM-4.5V 106B, InternVL3 78B) and proprietary models (Gemini 2.5 Pro, GPT-5 with high reasoning). Results reveal strikingly low performance (26.1–54.2%), underscoring the challenge of multimodal scientific reasoning and motivating progress towards trustworthy scientific assistants.
@@ 1 Introduction
Recent advances in Large Multimodal Models (LMMs) have sparked growing interest in their potential to serve as intelligent assistants for scientific research (Eger et al., 2025), supporting tasks such as figure and chart interpretation (Roberts et al., 2024; Tanasă & Oprea, 2025; Wu et al., 2024), paper summarization (Tan et al., 2025; Saxena et al., 2025; Yu et al., 2025), and error detection (Yang et al., 2025b; Miyai et al., 2024; Alsaif et al., 2024). Yet, a central open question remains: can LMMs truly reason over the complex multimodal structure of scientific documents?
@@
In this paper, we introduce PRISMM-Bench, a Peer-Review-sourced Inconsistency Set for Multimodal Models. Unlike synthetic datasets, PRISMM-Bench captures inconsistencies explicitly flagged by human reviewers in scientific papers on OpenReview. Through a multi-stage pipeline combining large-scale review mining, LLM-assisted filtering, and rigorous human verification, we curate a dataset of 262 inconsistencies across 242 papers submitted to ICLR 2025, spanning 13 categories of visual-textual and inter-visual mismatches. PRISMM-Bench provides a principled resource for evaluating and improving LMMs, grounded in the real challenges of understanding and verifying scientific papers. Grounded on these recent reviews, we minimize data contamination risk and demonstrate a pipeline with the potential to construct a continuously updated live benchmark. Building on this inconsistency set, we design a benchmark suite of three multiple-choice question (MCQ) tasks: 1) Inconsistency Identification - detect what the inconsistency is; 2) Inconsistency Remedy - determine how to fix the inconsistency; and 3) Inconsistency Pair Match - identify which two elements conflict. Together, these tasks form a tiered framework that evaluates not only a model’s ability to spot inconsistencies but also its capacity to propose remedy, and reason over relationships between different modality components.
@@
In MCQ evaluation, a key challenge is models’ tendency to rely on linguistic biases in answer choices. Prior work has shown that LLMs often exploit choice-only shortcuts, achieving non-trivial accuracy without reading the question (Chandak et al., 2025; Turner & Kurzeja, 2025; Balepur et al., 2024; Chizhov et al., 2025), and similar effects appear in multimodal MCQs (Chandak et al., 2025). To address this, we propose a novel structured JSON-based answer representation that de-emphasizes stylistic cues and minimizes spurious correlations. Inspired by  Das et al. (2014) and  Banarescu et al. (2013), our design converts free-form natural language into uniform structured representations that reduce model sensitivity to surface-level patterns. Our user study confirms that this approach suppresses linguistic shortcuts and better aligns models with human reasoning.
@@
We benchmark 21 state-of-the-art LMMs, spanning large open-weight models such as GLM-4.5V 106B (Hong et al., 2025) and InternVL3 78B (Zhu et al., 2025), as well as leading proprietary models including Gemini 2.5 Pro (Comanici et al., 2025) and GPT-5 (OpenAI, 2025). Results show that while large open-weight models achieve around 40% accuracy, even the strongest proprietary models reach just 54.2%, underscoring difficulty of the benchmark and limitations of current LMMs.
@@
Our contributions are fourfold: (1) We propose a reviewer-sourced dataset of real multimodal inconsistencies in scientific papers, spanning diverse categories and grounded in peer review. (2) We construct a benchmark suite of three tasks probing detection, correction, and relational reasoning over these inconsistencies. (3) We are the first to propose JSON-based debiasing method for MCQ, converting free-form responses into uniform structured representations. (4) We evaluate 21 state-of-the-art LMMs, exposing their current limitations in detecting, understanding and correcting inconsistencies in scientific papers. The dataset and code for creating the benchmark and evaluating LMMs are publicly available.
@@ 2 Related Work
Large Multimodal Models (LMMs). Large Multimodal Models (LMMs) pair a vision encoder with a large language model, enabling open-ended reasoning across tasks such as image captioning, VQA, document understanding, and chart interpretation. Early approaches like BLIP-2 (Li et al., 2023) and InstructBLIP (Dai et al., 2023) introduced instruction tuning on pre-trained vision-language models, while the LLaVA series (Liu et al., 2023b; a; Li et al., 2024a) advanced perception and reasoning via large-scale visual instruction tuning. Several recent studies (llava_icl; gavrikov2024vision; lin2024comparison; huang2024conme; glov; Hansen et al., 2025; Mei et al., 2025) have advanced these models by introducing improved training and adaptation strategies. Recent models extend these capabilities: Qwen-2.5 VL (Bai et al., 2025) offers precise object localization, dynamic resolution, and agentic tool execution; InternVL3 (Zhu et al., 2025) improves perception and reasoning through domain-specific pretraining on 3D scenes, GUIs, and video; Gemma 3 (Team et al., 2025), Ovis 2 (Lu et al., 2025), and GLM 4.5V (Hong et al., 2025) demonstrate strong performance across diverse multimodal benchmarks. High-resolution variants such as InternLM XComposer 2.5 (Zhang et al., 2024a) and VILA HD 4K (Shi et al., 2025), enable detailed perception and document processing. Proprietary models like GPT-5 (OpenAI, 2025) and Gemini 2.5 Pro (Comanici et al., 2025) set the state-of-the-art on complex multimodal tasks through large-scale training and enhanced reasoning.
@@
Language Biases in Evaluation Benchmarks. Multiple-choice evaluation is prone to linguistic biases, where models exploit surface-level patterns in answer options rather than reasoning over content. Prior studies show LLMs can achieve high accuracy even without the question, such as in TruthfulQA (Turner & Kurzeja, 2025), HellaSwag (Zellers et al., 2019), and ARC (Balepur et al., 2024). For example, Chandak et al. (2025) report 83% accuracy on TruthfulQA v2 using answer choices alone, with shortcut rates above 70% on HellaSwag. The recent trend of generating distractors with LLMs (e.g., in MMLU-Pro; Wang et al. (2024a)) can even exacerbate these artifacts.
@@ 3.1 Collection of Reviewer-Flagged Inconsistencies
To build a benchmark of realistic and authentic inconsistencies, we sourced cases flagged by reviewers on OpenReview (ope, b), where comments often highlight discrepancies between textual content and visual or mathematical components, including figures, tables, and equations.
Review Sourcing Strategy. We collected reviews from ICLR 2025 submissions via the OpenReview API v2 (ope, c). To maximize the likelihood that flagged inconsistencies persisted in the final public PDFs, we restricted to rejected or withdrawn submissions without rebuttals. [^1] Our earlier exploration of review sourcing revealed that most reviewer-flagged inconsistencies were resolved during rebuttal and did not persist in the final versions, motivating the current refined sourcing strategy. This yielded 12,366 reviews (details in App. D.1).
LLM Review Filtering. As manual screening for all reviews was infeasible, we employed an LLM for review filtering. Specifically, we used Mistral Nemo 2407 (Mistral, 2024) with low temperature settings to summarize reviews and identify potential inconsistency mentions, resulting in a curated set of 5,258 potential inconsistencies spanning 3,206 reviews (prompt details in App. F.1).
@@
This process produced 262 validated inconsistencies across 242 ICLR submissions. We identified 13 categories of inconsistencies based on the elements involved (distribution shown in Fig. 9). The most common cases were figure–text mismatches (24.4%) and intra-figure inconsistencies (24.0%).
@@ 3.2 Generation of Benchmark Tasks
Inconsistency Pair Match (Match). This task is built on a subset of inconsistencies that involve two distinct visual elements within a paper (135 samples). Given one element as context, the model must select its inconsistency counterpart from four options. By restricting the task to visual-visual mismatches, we specifically assess a model’s ability to detect representation errors without relying on textual cues, simulating the common peer-review challenge of cross-checking figures and tables for consistency.
More details about the task generation process are provided in App. D.4. We provide qualitative examples of the three tasks in App. A and a dataset viewing tool in the supplementary materials.
@@ 4.1 Experimental Setup
Model Selection. We evaluate 21 LMMs spanning a diverse range of architectures: 16 open-weight models of varying scales, two specialized high-resolution models, and three proprietary models. Selection was guided by performance on the Open VLM Leaderboard (ope, a) and the availability of the latest model versions.
@@ 4.2 Main Results
Gemma 3 4B 4.0 25.6 25.2 38.5 24.4 24.0 25.6 25.6 26.1
LLaVA OV 7B 7.0 29.4 26.3 23.7 32.4 27.9 29.0 27.1 28.3
Ovis2 8B 8.0 33.2 27.5 17.0 36.6 23.7 32.8 26.3 29.0
Qwen 2.5 VL 7B 7.0 30.9 28.2 59.3 29.0 29.8 26.7 29.0 31.3
InternVL3 8B 8.0 35.1 28.2 54.8 28.6 26.7 32.4 33.6 32.7
@@
Gemma 3 27B 27.0 34.4 30.5 64.4 27.9 26.0 30.5 26.7 32.1
Gemma 3 12B 12.0 32.1 30.5 64.4 32.4 26.0 30.5 26.7 32.5
Qwen 2.5 VL 32B 32.0 42.4 35.9 45.9 36.6 34.4 33.2 27.9 35.9
Ovis2 34B 34.0 49.6 40.5 33.3 40.1 35.1 33.2 32.1 38.0
InternVL3 38B 38.0 47.3 37.0 57.8 39.7 32.4 34.0 29.4 38.3
@@
LLaVA OV 72B 72.0 36.3 30.5 28.1 32.1 27.9 32.4 27.9 30.9
Qwen 2.5 VL 72B 72.0 50.8 34.7 26.7 41.6 29.0 36.3 23.3 35.2
InternVL3 78B 78.0 49.6 37.8 45.9 38.9 33.6 36.6 32.1 38.3
@@
Performance Landscape. Proprietary models substantially outperform their open-weight counterparts. Gemini 2.5 Pro and GPT-5 (high reasoning) reach the highest average performance of 54.2%. By contrast, the best open-weight model GLM 4.5V 106B achieves 41.9%, matching GPT-5 (minimal reasoning) but trailing its high-reasoning variant by 12.3 points. These results underscore the difficulty of the benchmark: even the best-performing models remain far from the reliability required of automated scientific assistants.
@@
Task formulation also plays a critical role. Remedy scores are consistently lower than Ident, showing that proposing corrections requires deeper reasoning than only detection. Performance on Match varies widely across models: Gemma 3 12B achieves 64.4%, rivaling proprietary models, whereas much larger models such as InternVL3 78B lag behind at 45.9%. These results suggest that architectural design, not just scale, is critical for relational reasoning.
Model Characteristics. Reasoning-enabled models show benefit. For example, InternVL3.5 8B outperforms its non-reasoning predecessor InternVL3 8B by over 5 percentage points, achieving accuracy comparable to models with nearly nine times more parameters. Turning off chain-of-thought reduces accuracy by up to 19 points (cf. App. C.2), directly confirming the contribution of reasoning. In contrast, high-resolution specialists (VILA HD 4K 8B, InternLM XC 2.5 7B) show little advantage in extended-context settings. More broadly, our results challenge the “bigger is better” paradigm: scaling up parameter counts alone does not guarantee higher performance, with diminishing returns observed from medium- to large-scale models.

### v3

@@ Abstract
Large Multimodal Models (LMMs) are increasingly applied to scientific research, yet it remains unclear whether they can reliably understand and reason over the multimodal complexity of papers. A central challenge lies in detecting and resolving inconsistencies across text, figures, tables, and equations, issues that are often subtle, domain-specific, and ultimately undermine clarity, reproducibility, and trust. Existing benchmarks overlook this issue, either isolating single modalities or relying on synthetic errors that fail to capture real-world complexity. We introduce PRISMM-Bench (Peer-Review-sourced Inconsistency Set for Multimodal Models), the first benchmark grounded in real reviewer-flagged inconsistencies in scientific papers. Through a multi-stage pipeline of review mining, LLM-assisted filtering and human verification, we curate 384 inconsistencies from 353 papers. Based on this set, we design three tasks, namely inconsistency identification, remedy and pair matching, which assess a model’s capacity to detect, correct, and reason over inconsistencies across different modalities. Furthermore, to address the notorious problem of choice-only shortcuts in multiple-choice evaluation, where models exploit answer patterns without truly understanding the question, we further introduce structured JSON-based answer representations that minimize linguistic biases by reducing reliance on superficial stylistic cues. We benchmark 21 leading LMMs, including large open-weight models (GLM-4.5V 106B, InternVL3 78B) and proprietary models (Gemini 2.5 Pro, GPT-5 with high reasoning). Results reveal strikingly low performance (27.8-53.9%), underscoring the challenge of multimodal scientific reasoning and motivating progress towards trustworthy scientific assistants.
@@ 1 Introduction
Recent advances in Large Multimodal Models (LMMs) have sparked growing interest in their potential to serve as intelligent assistants for scientific research (Eger et al., 2025), supporting tasks such as figure and chart interpretation (Roberts et al., 2024; Tanasă and Oprea, 2025; Wu et al., 2024), paper summarization (Tan et al., 2025; Saxena et al., 2025; Yu et al., 2025), and error detection (Yang et al., 2025b; Miyai et al., 2024; Alsaif et al., 2024). Yet, a central open question remains: can LMMs truly reason over the complex multimodal structure of scientific documents?
@@
In this paper, we introduce PRISMM-Bench, a Peer-Review-sourced Inconsistency Set for Multimodal Models. Unlike synthetic datasets, PRISMM-Bench captures inconsistencies explicitly flagged by human reviewers in scientific papers on OpenReview. Through a multi-stage pipeline combining large-scale review mining, LLM-assisted filtering, and rigorous human verification, we curate a dataset of 384 inconsistencies across 353 papers submitted to ICLR 2024 & 2025, spanning 15 categories of visual-textual and inter-visual mismatches. PRISMM-Bench provides a principled resource for evaluating and improving LMMs, grounded in the real challenges of understanding and verifying scientific papers. Grounded on these recent reviews, we minimize data contamination risk and demonstrate a pipeline with the potential to construct a continuously updated live benchmark. Building on this inconsistency set, we design a benchmark suite of three multiple-choice question (MCQ) tasks: 1) Inconsistency Identification - detect what the inconsistency is; 2) Inconsistency Remedy - determine how to fix the inconsistency; and 3) Inconsistency Pair Match - identify which two elements conflict. Together, these tasks form a tiered framework that evaluates not only a model’s ability to spot inconsistencies but also its capacity to propose remedy, and reason over relationships between different modality components.
@@
In MCQ evaluation, a key challenge is models’ tendency to rely on linguistic biases in answer choices. Prior work has shown that LLMs often exploit choice-only shortcuts, achieving non-trivial accuracy without reading the question (Chandak et al., 2025; Turner and Kurzeja, 2025; Balepur et al., 2024; Chizhov et al., 2025), and similar effects appear in multimodal MCQs (Chandak et al., 2025). To address this, we propose a novel structured JSON-based answer representation that de-emphasizes stylistic cues and minimizes spurious correlations. Inspired by  Das et al. (2014) and  Banarescu et al. (2013), our design converts free-form natural language into uniform structured representations that reduce model sensitivity to surface-level patterns. Our user study confirms that this approach suppresses linguistic shortcuts and better aligns models with human reasoning.
@@
We benchmark 21 state-of-the-art LMMs, spanning large open-weight models such as GLM-4.5V 106B (Hong et al., 2025) and InternVL3 78B (Zhu et al., 2025), as well as leading proprietary models including Gemini 2.5 Pro (Comanici et al., 2025) and GPT-5 (OpenAI, 2025). Results show that while large open-weight models achieve around 40% accuracy, even the strongest proprietary models reach just 53.9%, underscoring difficulty of the benchmark and limitations of current LMMs.
@@
Our contributions are fourfold: (1) We propose a reviewer-sourced dataset of real multimodal inconsistencies in scientific papers, spanning diverse categories and grounded in peer review. (2) We construct a benchmark suite of three tasks probing detection, correction, and relational reasoning over these inconsistencies. (3) We are the first to propose JSON-based debiasing method for MCQ, converting free-form responses into uniform structured representations. (4) We evaluate 21 state-of-the-art LMMs, exposing their current limitations in detecting, understanding and correcting inconsistencies in scientific papers.
@@ 2 Related Work
Large Multimodal Models (LMMs). Large Multimodal Models (LMMs) pair a vision encoder with a large language model, enabling open-ended reasoning across tasks such as image captioning, VQA, document understanding, and chart interpretation. Early approaches like BLIP-2 (Li et al., 2023) and InstructBLIP (Dai et al., 2023) introduced instruction tuning on pre-trained vision-language models, while the LLaVA series (Liu et al., 2023b; a; Li et al., 2024a) advanced perception and reasoning via large-scale visual instruction tuning. Several recent studies (Doveh et al., 2024; Gavrikov et al., 2024; Lin et al., 2024; Huang et al., 2024; Mirza et al., 2025; Hansen et al., 2025; Mei et al., 2025) have advanced these models by introducing improved training and adaptation strategies. Recent models extend these capabilities: Qwen-2.5 VL (Bai et al., 2025) offers precise object localization, dynamic resolution, and agentic tool execution; InternVL3 (Zhu et al., 2025) improves perception and reasoning through domain-specific pretraining on 3D scenes, GUIs, and video; Gemma 3 (Team et al., 2025), Ovis 2 (Lu et al., 2025), and GLM 4.5V (Hong et al., 2025) demonstrate strong performance across diverse multimodal benchmarks. High-resolution variants such as InternLM XComposer 2.5 (Zhang et al., 2024a) and VILA HD 4K (Shi et al., 2025), enable detailed perception and document processing. Proprietary models like GPT-5 (OpenAI, 2025) and Gemini 2.5 Pro (Comanici et al., 2025) set the state-of-the-art on complex multimodal tasks through large-scale training and enhanced reasoning.
@@
Language Biases in Evaluation Benchmarks. Multiple-choice evaluation is prone to linguistic biases, where models exploit surface-level patterns in answer options rather than reasoning over content. Prior studies show LLMs can achieve high accuracy even without the question, such as in TruthfulQA (Turner and Kurzeja, 2025), HellaSwag (Zellers et al., 2019), and ARC (Balepur et al., 2024). For example, Chandak et al. (2025) report 83% accuracy on TruthfulQA v2 using answer choices alone, with shortcut rates above 70% on HellaSwag. The recent trend of generating distractors with LLMs (e.g., in MMLU-Pro; Wang et al. (2024a)) can even exacerbate these artifacts.
@@ 3.1 Collection of Reviewer-Flagged Inconsistencies
To build a benchmark of realistic and authentic inconsistencies, we sourced cases flagged by reviewers on OpenReview (55), where comments often highlight discrepancies between textual content and visual or mathematical components, including figures, tables, and equations.
Review Sourcing Strategy. We collected reviews from ICLR 2024 & 2025 submissions via the OpenReview API v2 (54). To maximize the likelihood that flagged inconsistencies persisted in the final public PDFs, we restricted to rejected or withdrawn submissions without rebuttals. [^1] Our earlier exploration of review sourcing revealed that most reviewer-flagged inconsistencies were resolved during rebuttal and did not persist in the final versions, motivating the current refined sourcing strategy. This yielded 18,009 reviews (details in App. D.1).
LLM Review Filtering. As manual screening for all reviews was infeasible, we employed an LLM for review filtering. Specifically, we used Mistral Nemo 2407 (Mistral, 2024) with low temperature settings to summarize reviews and identify potential inconsistency mentions, resulting in a curated set of 6,056 potential inconsistencies spanning 2,458 reviews (prompt details in App. F.1).
@@
This process produced 384 validated inconsistencies across 353 ICLR submissions. We identified 15 categories of inconsistencies based on the elements involved (distribution shown in Fig. 9). The most common cases were intra-figure inconsistencies (23.7%) and figure-text mismatches (21.9%).
@@ 3.2 Generation of Benchmark Tasks
Inconsistency Pair Match (Match). This task is built on a subset of inconsistencies that involve two distinct visual elements within a paper (192 samples). Given one element as context, the model must select its inconsistency counterpart from four options. By restricting the task to visual-visual mismatches, we specifically assess a model’s ability to detect representation errors without relying on textual cues, simulating the common peer-review challenge of cross-checking figures and tables for consistency.
More details about the task generation process and a validation of the multiple-choice format against open-ended evaluation are provided in App. D.4 and App. C.3, respectively. We provide qualitative examples of the three tasks in App. A and a dataset viewing tool in the supplementary materials.
@@ 4.1 Experimental Setup
Model Selection. We evaluate 21 LMMs spanning a diverse range of architectures: 16 open-weight models of varying scales, two specialized high-resolution models, and three proprietary models. Selection was guided by performance on the Open VLM Leaderboard (52) and the availability of the latest model versions.
@@ 4.2 Main Results
Gemma 3 4B 4.0 27.9 29.9 39.6 25.0 24.7 26.6 27.1 27.8
LLaVA OV 7B 7.0 30.5 28.4 29.7 32.0 28.4 28.1 27.9 29.2
Ovis2 8B 8.0 35.4 29.4 22.4 34.4 27.3 31.8 28.1 30.4
Qwen 2.5 VL 7B 7.0 32.8 31.3 58.9 29.9 29.7 26.8 27.1 31.9
InternVL3 8B 8.0 36.5 29.4 56.3 28.6 27.6 30.7 31.8 32.7
@@
Gemma 3 27B 27.0 36.2 32.8 59.9 30.7 28.6 31.0 27.3 33.3
Gemma 3 12B 12.0 33.9 30.5 63.5 30.7 25.8 30.7 30.5 32.9
Qwen 2.5 VL 32B 32.0 42.4 37.0 45.8 37.2 34.6 38.3 27.9 37.0
Ovis2 34B 34.0 50.0 41.1 37.0 40.6 36.2 33.3 31.8 38.7
InternVL3 38B 38.0 46.6 38.5 56.8 40.6 35.7 37.0 32.0 39.8
@@
LLaVA OV 72B 72.0 35.4 30.5 28.1 32.3 28.4 31.5 26.0 30.5
Qwen 2.5 VL 72B 72.0 49.7 37.2 32.8 44.0 33.3 35.4 25.3 37.1
InternVL3 78B 78.0 49.5 39.3 45.3 39.3 33.9 35.9 30.5 38.6
@@
Performance Landscape. Proprietary models substantially outperform their open-weight counterparts. GPT-5 (high reasoning) reached the highest average performance of 53.9%. By contrast, the best open-weight model GLM 4.5V 106B achieves 42.5%, matching GPT-5 (minimal reasoning) but trailing its high-reasoning variant by 11.4 points. These results underscore the difficulty of the benchmark: even the best-performing models remain far from the reliability required of automated scientific assistants.
@@
Task formulation also plays a critical role. Remedy scores are consistently lower than Ident, showing that proposing corrections requires deeper reasoning than only detection. Performance on Match varies widely across models: Gemma 3 12B achieves 63.5%, rivaling proprietary models, whereas much larger models such as InternVL3 78B lag behind at 45.3%. These results suggest that architectural design, not just scale, is critical for relational reasoning.
Model Characteristics. Reasoning-enabled models show benefit. For example, InternVL3.5 8B outperforms its non-reasoning predecessor InternVL3 8B by over 5 percentage points, achieving accuracy comparable to models with nearly nine times more parameters. Turning off chain-of-thought reduces accuracy by up to 14 points (cf. App. C.2), directly confirming the contribution of reasoning. In contrast, high-resolution specialists (VILA HD 4K 8B, InternLM XC 2.5 7B) show little advantage in extended-context settings. More broadly, our results challenge the “bigger is better” paradigm: scaling up parameter counts alone does not guarantee higher performance, with diminishing returns observed from medium- to large-scale models.
@@
4.3 Analysis of Chain-of-Thought Reasoning
Reasoning variants perform better than non-reasoning counterparts and reach results comparable to much larger models. For instance, InternVL3.5 8B achieves an average of 37.7%, rivaling large open-weight models and surpassing several 72B non-reasoning models. We therefore study how reasoning-enabled models leverage chain-of-thought (CoT) to improve performance on PRISMM-Bench. To do so, we re-evaluated a selection of reasoning models on the Ident task with Focused Context with reasoning turned off and compared the performance.
Ablation Results. Disabling reasoning reveals the critical role of CoT in detecting subtle inconsistencies. For example, GLM 4.5V drops from 51.8% to 43.2% (-16.6%), InternVL3.5 8B from 49.5% to 40.6% (-18.0%), and InternVL3.5 38B suffers the largest decline, from 54.4% to 40.4% (-25.7%).
Why Reasoning Helps. To understand these performance differences, we focused on inconsistencies where InternVL3.5 38B succeeded with reasoning but failed without. We identified three consistent patterns: (1) Structured input handling: Reasoning-enabled models interpret the JSON-formatted options in natural language, clarifying subtle distinctions without exploring linguistic biases (cf. low without-context performance for reasoning models in Sec. 4.4). (2) Cross-modal grounding: CoT traces show models explicitly reasoning over both text and visuals, breaking complex information into smaller units and reusing them later in the reasoning chain. (3) Concept linking: Reasoning enables models to connect fine-grained context with domain knowledge and abstract concepts, allowing stronger logical inference beyond surface pattern recognition. We provide a detailed case study illustrating these effects in App. C.2.